# A Formalism for the Transport and Matching of Coupled Beams in Accelerators

### v1

@@ Abstract
Understanding transverse coupling dynamics is crucial for beam physics, accelerator design, and operations. Currently, most accelerators are designed for uncoupled beams, and coupling is treated as an error or perturbation. Many transverse ($x$-$y$) coupling parametrizations exist: Edward-Teng, Mais-Ripken, Levedev-Bogacz, and others. Here, we present an explicit and complete formalism for transporting coupled beam optics functions based on Mais-Ripken and Lebedev-Bogacz formalism. The formalism allows for matching generally coupled beam optics functions but applies to uncoupled optics as well. A complete transformation method for coupled optics provides easy matching routines that can be added to known beam optics codes that lack this feature. For fully coupled lattices, we present methods for extracting eigenmode emittances and other beam parameters from observables that can be measured, which is essential to diagnose and characterize the beam in a real machine.
@@ I Introduction
Conventional accelerators are typically based on uncoupled beam dynamics. Courant and Snyder developed the uncoupled beam parametrization and defined the well-known beam optics functions, also called Twiss parameters [1]. Uncoupled beam dynamics have been studied extensively, and important beam properties have been determined and characterized, such as beam stability and perturbations [2, 3]. Single-particle dynamics shed light on beam resonances and the importance of machine operating points. Uncoupled single-particle dynamics are relatively easy to deal with due to the separation of transverse planes, where the primary source of coupling is magnet alignment errors in a linear machine. Usually, the weak coupling created in the machine is corrected using skew quadrupoles. However, coupling never disappears in the lattice since higher-order magnets are used for various corrections, and fringe fields always introduce a source of coupling to the system, whether linear or nonlinear. Recently, there has been a growing interest in the use of coupled beam optics and coupled lattices as options that can lead to higher luminosities and beam brightness.
Single-particle coupled parametrizations have already been introduced. First, Mais-Ripken introduced their coupled parametrization as an extension of the Courant-Snyder parametrization [4]. Edwards-Teng introduced a general decoupled parametrization through a transformation that removes the coupling and considers the optics as in the uncoupled case of Courant-Snyder parametrization [5]. Lebedev-Bogacz introduced an extension of Mais-Ripken parametrization by parameterizing the transfer map and writing down the eigenvectors [6]. Most of these parametrizations are presented and summarized in this review on coupling [7]. New coupled parametrization approaches have been published recently, shedding more light on beam coupling [8]. In [8], they also show the tune map, which suggests that the resonance diagram will be surfaces rather than resonance lines, which is an interesting phenomenon. Recent advancements in 4D phase space reconstruction make coupled lattices possible [9, 10]. Turn-by-turn data can also give clues about the coupled parameters, as shown in [11]. There is also interesting work suggesting that specific strong correlations between $x$ and $y$ planes can lead to better space-charge performance and even self-consistent beam distributions [12, 13, 14, 15]. A fully coupled beam with nonzero angular momentum creates a delta function correlation between the $x$$y^{\prime}$ and $y$$x^{\prime}$ planes, which leads to intrinsic flatness with a round beam cross-section, called circular modes. Circular mode beams inherit strong $x$$y$ plane coupling and intrinsic flatness through Derbenev’s Adapter, for example, which is presented in the Applications section. Danilov distribution is a class of coupled beam distributions that are self-consistent under space charge forces, similar to Kapchinskij-Vladimirskij (KV) distribution [16]. Furthermore, it is proposed that synchrotron sources can benefit from round beams rather than flat beams, especially for low-emittance machines [17], where manipulation of the beam shape can be achieved through managing coupling scales and strengths in a lattice for the desired beam projections.
@@
The most important challenge in designing a linear coupled lattice is the behavior of the optics functions. This paper follows Mais-Ripken and Lebedev-Bogacz parametrizations of the optics functions, introducing additional optics functions. Their periodic behavior is essential in designing coupled lattices. First, we review the Courant-Snyder parametrization and introduce the generating vector formalism and how the generating vectors transform, reproducing the known transfer matrix formalism. Secondly, we extend the generating vector formalism to the coupled case using a method similar to Mais-Ripken parametrization. We also make a connection between the Mais-Ripken and Lebedev-Bogacz parameters. We then write the explicit transfer matrix method for the coupled optics functions of eigenmodes $1$ and $2$, similar to the Courant-Snyder transfer matrix. The transfer of optics functions is derived from the generating vectors. Using the Lebedev-Bogacz parametrization, we show that the generating vectors also decouple using the decoupling matrix as in the Edwards-Teng parametrization. Following the description of the formalism, we present some application cases of the transformation and compare them with existing beam optics codes such as MAD-X and OptiMX. An important application is the periodic matching of coupled optics functions. Popular optics codes such as MAD-X and OptiMX come with the calculation of coupled optics functions for a given lattice and initial uncoupled functions. Both MAD-X and OptiMX provide matching routines that can compute periodic solutions for uncoupled optics functions using Courant-Snyder parametrization or Edward-Teng parametrization. However, there is no matching routine for coupled optics functions. The formulation in this paper was used to construct a matching routine for generally coupled beams, which could be adapted for popular optics and design codes. Finally, using this formalism, we developed a fully coupled and periodic ring design.
@@ III Coupled Beam Optics parametrizations
$\vec{z}=\frac{1}{2}\sqrt{\epsilon_{1}}e^{-i\psi_{1}}\vec{v}_{1}+\frac{1}{2}\sqrt{\epsilon_{2}}e^{-i\psi_{2}}\vec{v}_{2}+\frac{1}{2}\sqrt{\epsilon_{1}}e^{i\psi_{1}}\vec{v}_{1}^{*}+\frac{1}{2}\sqrt{\epsilon_{2}}e^{i\psi_{2}}\vec{v}_{2}^{*}.$ (12)
Here, $\epsilon_{1,2}$ are the invariants (emittances) of the system, $\psi_{1,2}$ are the initial betatron phases, $\vec{v}_{1,2}$ are the eigenvectors. The eigenvectors are given by:
@@
where $\nu_{1,2}$ are defined as the phases of coupling and $u$ the coupling strength parameter. The eigenvectors are normalized with respect to the symplectic unit matrix, $\vec{v}_{1,2}^{\dagger}S\vec{v}_{1,2}=-2i$. Application of the transfer matrix to the eigenvectors yields the phase advances as eigenvalues. Therefore, this leads to an equation similar to Eq. (7):
@@
$\begin{split}\vec{z}_{1}&=\frac{1}{2}(\vec{v}_{1}+\vec{v}_{1}^{*})\cos\mu_{1}-\frac{i}{2}(\vec{v}_{1}-\vec{v}_{1}^{*})\sin\mu_{1},\\ \vec{z}_{2}&=\frac{1}{2}(\vec{v}_{1}+\vec{v}_{1}^{*})\sin\mu_{1}+\frac{i}{2}(\vec{v}_{1}-\vec{v}_{1}^{*})\cos\mu_{1},\\ \vec{z}_{3}&=\frac{1}{2}(\vec{v}_{2}+\vec{v}_{2}^{*})\cos\mu_{2}-\frac{i}{2}(\vec{v}_{2}-\vec{v}_{2}^{*})\sin\mu_{2},\\ \vec{z}_{4}&=\frac{1}{2}(\vec{v}_{2}+\vec{v}_{2}^{*})\sin\mu_{2}+\frac{i}{2}(\vec{v}_{2}-\vec{v}_{2}^{*})\cos\mu_{2}.\\ \end{split}$ (15)
@@
$\begin{split}\overline{\mu}_{1}&=\mu_{1}-\arctan\left(\frac{(1-u)}{\alpha_{1x}}\right),\quad\overline{\mu}_{2}=\mu_{2}-\arctan\left(\frac{1-u}{\alpha_{2y}}\right),\\ \overline{(\mu_{1}-\nu_{1})}&=(\mu_{1}-\nu_{1})-\arctan\left(\frac{u}{\alpha_{1y}}\right),\quad\overline{(\mu_{2}-\nu_{2})}=(\mu_{2}-\nu_{2})-\arctan\left(\frac{u}{\alpha_{2x}}\right).\\ \end{split}$ (19)
@@
$\begin{split}\vec{z}_{1}&=\begin{pmatrix}\sqrt{\beta_{1x}}\cos\mu_{1}\\ -\sqrt{\gamma_{1x}}\cos\overline{\mu}_{1}\\ \sqrt{\beta_{1y}}\cos(\mu_{1}-\nu_{1})\\ -\sqrt{\gamma_{1y}}\cos(\overline{\mu_{1}-\nu_{1}})\end{pmatrix},\quad\vec{z}_{2}=\begin{pmatrix}\sqrt{\beta_{1x}}\sin\mu_{1}\\ -\sqrt{\gamma_{1x}}\sin\overline{\mu}_{1}\\ \sqrt{\beta_{1y}}\sin(\mu_{1}-\nu_{1})\\ -\sqrt{\gamma_{1y}}\sin\overline{(\mu_{1}-\nu_{1})}\end{pmatrix},\\ \vec{z}_{3}&=\begin{pmatrix}\sqrt{\beta_{2x}}\cos(\mu_{2}-\nu_{2})\\ -\sqrt{\gamma_{2x}}\cos\overline{(\mu_{2}-\nu_{2})}\\ \sqrt{\beta_{2y}}\cos\mu_{2}\\ -\sqrt{\gamma_{2y}}\cos\overline{\mu}_{2}\end{pmatrix},\quad\vec{z}_{4}=\begin{pmatrix}\sqrt{\beta_{2x}}\sin(\mu_{2}-\nu_{2})\\ -\sqrt{\gamma_{2x}}\sin\overline{(\mu_{2}-\nu_{2})}\\ \sqrt{\beta_{2y}}\sin\mu_{2}\\ -\sqrt{\gamma_{2y}}\sin\overline{\mu}_{2}\end{pmatrix}.\end{split}$ (20)
@@ III.1 Relation to Edwards-Teng parametrization
$\begin{split}\begin{pmatrix}\sqrt{\beta_{1}}\\ -\frac{i+\alpha_{1}}{\sqrt{\beta_{1}}}\\ 0\\ 0\end{pmatrix}=\begin{pmatrix}\cos\phi&0&-d\sin\phi&b\sin\phi\\ 0&\cos\phi&c\sin\phi&-a\sin\phi\\ a\sin\phi&b\sin\phi&\cos\phi&0\\ c\sin\phi&d\sin\phi&0&\cos\phi\end{pmatrix}\cdot\begin{pmatrix}\sqrt{\beta_{1x}}\\ -\frac{i(1-u)+\alpha_{1x}}{\sqrt{\beta_{1x}}}\\ \sqrt{\beta_{1y}}e^{i\nu_{1}}\\ -\frac{iu+\alpha_{1y}}{\sqrt{\beta_{1y}}}e^{i\nu_{1}}\end{pmatrix},\end{split}$ (32)
@@
$\begin{split}\vec{\tilde{z}}_{1}&=\tilde{R}\vec{z}_{1}=\frac{1}{2}(\tilde{R}\vec{v}_{1}+\tilde{R}\vec{v}_{1}^{*})\cos\mu_{1}-\frac{i}{2}(\tilde{R}\vec{v}_{1}-\tilde{R}\vec{v}_{1}^{*}),\\ &=\frac{1}{2}(\tilde{v}_{1}+\tilde{v}_{1}^{*})\cos\mu_{1}-\frac{i}{2}(\tilde{v}_{1}-\tilde{v}_{1}^{*})\sin\mu_{1},\\ \end{split}$ (34)
@@
$\begin{split}\tilde{z}_{1}&=\begin{pmatrix}\sqrt{\beta_{1}}\cos\mu_{1}\\ \frac{-\alpha_{1}\cos\mu_{1}-\sin\mu_{1}}{\sqrt{\beta_{1}}}\\ 0\\ 0\end{pmatrix},\quad\tilde{z}_{2}=\begin{pmatrix}\sqrt{\beta_{1}}\sin\mu_{1}\\ \frac{-\alpha_{1}\sin\mu_{1}+\cos\mu_{1}}{\sqrt{\beta_{1}}}\\ 0\\ 0\end{pmatrix}.\end{split}$ (35)
@@ III.2 Special Solutions
For small coupling to no coupling, this suggests $u=0$ leads to correct assignment of emittances with $\epsilon_{1}=\epsilon_{x}$ and $\epsilon_{2}=\epsilon_{y}$. The set of relationships shown in Eq. (39) gives matched solutions for coupled functions while maintaining the normalization conditions between the generating vectors. When $u\rightarrow 0$, the coupling vanishes and it removes the ambiguity of setting the coupled beta functions to zero, which causes singularity. This set of solutions is illustrated in Fig. 2.
@@ IV Applications: Lattices & Lattice Building Blocks
In this section, we apply this method to various examples of lattices and lattice building blocks. In order to write a compact matching and transport script, the generating functions are used, where the relations of optics functions to the generating functions are given in Eq. (23). It is important to check the normalization condition for the transport of the generating vectors.
@@ IV.1 Derbenev’s Adapter
Derbenev’s Adapter, known as a flat-to-round beam converter, is a triplet of skew quadrupoles that converts flat beams into coupled round beams with non-zero angular momentum known as circular modes [18, 12, 19]. This transformation creates a strong coupling between the $x$ and $y$ planes, which produces fully coupled optics. Fully coupled optics correspond to coupling strength parameter, $u=1/2$ [6]. Using the method developed here, we can match the three skew quadrupoles to perform a flat-to-round transformation. The matched results for an initial uncoupled beam with $\beta_{x,y}=5$ m are shown in Fig. 3. Derbenev adapter reduces the uncoupled functions by half, creating a strong $x-y$ coupling. Strong coupling is seen at $u=1/2$. After the transformation, the phases of coupling are $\pi/2$ and the $x-y$ cross section is round. Transformation of the optics functions from uncoupled to coupled optics can be done using Eq. (29):
@@ IV.2 Mode Flipping with Skew Quadrupoles
Flat beams are beams in which the emittance is much smaller in one plane than in the other, for example $\epsilon_{y}\ll\epsilon_{x}$, when the beam is flat in the $y$ direction. The beam sizes are usually larger in the $x$ dimension than in the $y$ dimension, mainly due to dispersion. Flatness is usually achieved for electrons in storage rings as a result of damping from synchotron radiation. The flatness can be exchanged between the $x$ and $y$ planes using a set of skew quadrupoles. As we have discussed $u=0$ corresponds to uncoupled motion. This suggests that mode labeling can be $\epsilon_{x}\rightarrow\epsilon_{1}$ and $\epsilon_{y}\rightarrow\epsilon_{2}$, for example. We also notice that $u=1$ leads to generating vectors with similar form as the case with $u=0$, but with mode flipping. Using the transformation of optics functions above, we can design a set of skew quadrupoles that achieve mode flipping by taking $u=0$ to $u=1$ as shown in Fig. 4. The process of mode flipping is simply to exchange the modes’ contributions to x and y projections: $\epsilon_{x}\rightarrow\epsilon_{1}\rightarrow\epsilon_{y}$ and $\epsilon_{y}\rightarrow\epsilon_{2}\rightarrow\epsilon_{x}$. This flips an initial flat beam in $y$ with $\epsilon_{y}\ll\epsilon_{x}$ into a flat beam in $x$ with $\epsilon_{x}\ll\epsilon_{y}$.
@@ IV.4 Solenoid-Based Cells
Solenoid focusing is mainly used at low energies; it is less effective and requires very high fields at high energies. Solenoids are used to propagate axisymmetric beams, to produce magnetized beams, and for beam cooling. Interestingly, solenoids can split the tunes of the coupled beams into mode $1$ and mode $2$ tunes. The splitting of tunes is illustrated in Fig. 6 for the same cell of two solenoids, one case with opposite polarity $K=-K$ and the other with the same polarity. In both cases, the propagation of beta functions is the same, whereas the effect on the phase advance is different; a splitting is observed. The solenoid cells were matched for $u=1/2$ and equal beta functions for a round beam. The propagation of phase advance in solenoid-type cells is given by
$\begin{split}\frac{d\mu_{1}}{ds}&=\frac{1-u}{\beta_{1x}}-\frac{R}{2}\sin\nu_{1},\\ \frac{d\mu_{2}}{ds}&=\frac{1-u}{\beta_{2y}}+\frac{R}{2}\sin\nu_{2}.\end{split}$ (43)
@@
Figure 6: Solenoid cells: Beta functions of opposite polarity of solenoid fields (top-left), same polarity (bottom-left). Phase advances of modes $1$ and $2$ of opposite polarity (top-right), same polarity (bottom-right), for $u=1/2$.

### v2

@@ Abstract
Understanding transverse coupling dynamics is crucial for beam physics, accelerator design, and operations. Currently, most accelerators are designed for uncoupled beams, and coupling is treated as an error or perturbation. Many transverse ($x$-$y$) coupling parametrizations exist: Edward-Teng, Mais-Ripken, Levedev-Bogacz, and others. Here, we present an explicit and complete formalism for transporting coupled beam optics functions based on Mais-Ripken and Lebedev-Bogacz parametrizations. The formalism allows for matching generally coupled beam optics functions but applies to uncoupled optics as well. A complete transformation method for coupled optics provides easy matching routines that can be added to known beam optics codes that lack this feature. For fully coupled lattices, we present methods for extracting eigenmode emittances and other beam parameters from observables that can be measured, which is essential to diagnose and characterize the beam in a real machine. We express the linear difference resonance in terms of coupled optics functions and relate it to the coupling strength parameter with explicit emittance exchange formalism realized from the generating functions discussed here.
@@ I Introduction
Conventional accelerators are typically based on uncoupled beam dynamics. Courant and Snyder developed the uncoupled beam parametrization and defined the well-known beam optics functions, also called Twiss parameters [1]. Uncoupled beam dynamics have been studied extensively, and important beam properties have been determined and characterized, such as beam stability and perturbations [2, 3]. Single-particle dynamics shed light on beam resonances and the importance of machine operating points. Uncoupled single-particle dynamics are relatively easy to deal with due to the separation of transverse planes, where the primary source of coupling is magnet alignment errors in a linear machine. Usually, the weak coupling created in the machine is corrected using skew quadrupoles. However, coupling never disappears in the lattice since higher-order magnets are used for various corrections, and fringe fields always introduce a source of coupling to the system, whether linear or nonlinear.
[…]
The introduction of linear coupling elements to the lattice adds new resonance conditions in addition to the integer resonances, namely sum and difference resonances originating from skew quadrupoles. The difference resonance is responsible for the emittance exchange mechanism in flat beams, where the emittance planes can be exchanged, resulting in a swap between $x$ and $y$ planes. Typically, a skew element is treated as a perturbation to the system, and the effect of the exchange rate is studied with respect to the resonance proximity parameter, $\Delta=Q_{x}-Q_{y}-\delta$, discussed in [18, 19]. The emittance exchange mechanism is described in terms of the resonance proximity parameter and the integrated skew strengths, where the invariant of the system is replaced by the sum of plane invariants. It is interesting to note that the stable difference resonance is studied from the perspective of uncoupled elements using perturbation theory. In this manuscript, we will introduce the emittance exchange mechanism for difference resonance from the perspective of coupled functions and parameters.
@@
The most important challenge in designing a linear coupled lattice is the behavior of the optics functions. This paper follows Mais-Ripken and Lebedev-Bogacz parametrizations of the optics functions. Their periodic behavior is essential in designing coupled lattices. First, we review the Courant-Snyder parametrization and introduce the generating vector formalism, explaining how the generating vectors transform to reproduce the known transfer matrix formalism. Secondly, we extend the generating vector formalism to the coupled case using a method similar to Mais-Ripken parametrization. We also make a connection between the Mais-Ripken and Lebedev-Bogacz parameters. We then write the explicit transfer matrix method for the coupled optics functions of eigenmodes $1$ and $2$, similar to the Courant-Snyder transfer matrix. The transfer of optics functions is derived from the generating vectors. Using the Lebedev-Bogacz parametrization, we show that the generating vectors also decouple using the decoupling matrix, similar to the Edwards-Teng parametrization. Following the description of the formalism, we present some application cases of the transformation and compare them with existing beam optics codes such as MAD-X and OptiMX. An important application is the periodic matching of coupled optics functions. Popular optics codes such as MAD-X and OptiMX come with the calculation of coupled optics functions for a given lattice and initial uncoupled functions. Both MAD-X and OptiMX provide matching routines that can compute periodic solutions for uncoupled optics functions using Courant-Snyder parametrization or Edward-Teng parametrization. However, there is no matching routine for coupled optics functions. The formulation in this paper was used to construct a matching routine for generally coupled beams, which could be adapted for popular optics and design codes. Using the formalism described here, we also apply it to the linear difference resonance and express the exchange mechanism from the perspective of coupled functions. Finally, using this formalism, we developed a fully coupled and periodic ring design.
@@ III Coupled Beam Optics parametrizations
$\vec{z}=\frac{1}{2}\sqrt{2J_{1}}e^{-i\psi_{1}}\vec{v}_{1}+\frac{1}{2}\sqrt{2J_{2}}e^{-i\psi_{2}}\vec{v}_{2}+\frac{1}{2}\sqrt{2J_{1}}e^{i\psi_{1}}\vec{v}_{1}^{*}+\frac{1}{2}\sqrt{2J_{2}}e^{i\psi_{2}}\vec{v}_{2}^{*}.$ (12)
Here, $J_{1,2}$ are the invariants of the system, $\psi_{1,2}$ are the initial betatron phases, $\vec{v}_{1,2}$ are the eigenvectors. The eigenvectors are given by:
@@
where $\nu_{1,2}$ are defined as the phases of coupling and $u$ the coupling strength parameter as defined in Appendix B. The coupling strength parameter indicates the weight of the mode contributions to the phase planes, while phases of coupling indicate how different modes project onto phase planes. The eigenvectors are normalized with respect to the symplectic unit matrix, $\vec{v}_{1,2}^{\dagger}S\vec{v}_{1,2}=-2i$. Application of the transfer matrix to the eigenvectors yields the phase advances as eigenvalues. Therefore, this leads to an equation similar to Eq. (7):
@@
[…]
@@
[…]
@@
[…]
@@ III.1 Relation to Edwards-Teng parametrization
[…]
@@
[…]
@@
[…]
@@ III.2 Special Solutions
For small coupling to no coupling, this suggests $u=0$ leads to correct assignment of emittances with $\epsilon_{1}=\epsilon_{x}$ and $\epsilon_{2}=\epsilon_{y}$. The set of relationships shown in Eq. (39) gives matched solutions for coupled functions while maintaining the normalization conditions between the generating vectors. When $u\rightarrow 0$, Eq. (13) do not fully reduce to the uncoupled eigenvectors, and further requirement of setting off-mode functions to zero is needed, which leads to singularity (division by zero). The special set of solutions for off-mode functions in Eq. (39) fixes this singularity for uncoupled lattice. This set of solutions is illustrated in Fig. 2.
@@ IV Applications: Lattices & Lattice Building Blocks
In this section, we apply this method to various examples of lattices and lattice building blocks. In order to write a compact matching and transport script, the generating functions are used, where the relations of optics functions to the generating functions are given in Eq. (23). It is important to check the normalization condition for the transport of the generating vectors which is done through the normalization condition. The python scripts that are used for matching and propagating of the coupled lattices is available on [20].
@@ IV.1 Derbenev’s Adapter
Derbenev’s Adapter, known as a flat-to-round beam converter, is a triplet of skew quadrupoles that converts flat beams into coupled round beams with non-zero angular momentum known as circular modes [21, 12, 22]. This transformation creates a strong coupling between the $x$ and $y$ planes, which produces fully coupled optics. Fully coupled optics correspond to coupling strength parameter, $u=1/2$ [6]. Using the method developed here, we can match the three skew quadrupoles to perform a flat-to-round transformation. The matched results for an initial uncoupled beam with $\beta_{x,y}=5$ m are shown in Fig. 3. Derbenev adapter reduces the uncoupled functions by half, creating a strong xyx-y coupling. Strong coupling is seen at $u=1/2$. After the transformation, the phases of coupling are $\pi/2$ and the xyx-y cross section is round. Transformation of the optics functions from uncoupled to coupled optics can be done using Eq. (29):
@@ IV.2 Mode Flipping with Skew Quadrupoles
Uncoupled flat beams are beams in which the emittance is much smaller in one plane than in the other, for example $\epsilon_{y}\ll\epsilon_{x}$, when the beam is flat in the $y$ direction. The beam sizes are usually larger in the $x$ dimension than in the $y$ dimension. Flatness is usually achieved for electrons in storage rings as a result of damping from synchotron radiation. The flatness can be exchanged between the $x$ and $y$ planes using a set of skew quadrupoles. As we have discussed $u=0$ corresponds to uncoupled motion. This suggests that mode labeling can be $\epsilon_{x}\rightarrow\epsilon_{1}$ and $\epsilon_{y}\rightarrow\epsilon_{2}$, for example. We also notice that $u=1$ leads to generating vectors with similar form as the case with $u=0$, but with mode flipping. Using the transformation of optics functions above, we can design a set of skew quadrupoles that achieve mode flipping by taking $u=0$ to $u=1$ as shown in Fig. 4. The process of mode flipping is simply to exchange the modes’ contributions to x and y projections: $\epsilon_{x}\rightarrow\epsilon_{1}\rightarrow\epsilon_{y}$ and $\epsilon_{y}\rightarrow\epsilon_{2}\rightarrow\epsilon_{x}$. This flips an initial uncoupled flat beam in $y$ with $\epsilon_{y}\ll\epsilon_{x}$ into an uncoupled flat beam in $x$ with $\epsilon_{x}\ll\epsilon_{y}$.
@@ IV.4 Solenoid-Based Cells
Solenoid focusing is mainly used at low energies; it is less effective and requires very high fields at high energies. Solenoids are used to propagate axisymmetric beams, to produce magnetized beams, and for beam cooling. Interestingly, solenoids can split the tunes of the coupled beams into mode $1$ and mode $2$ tunes. The splitting of tunes is illustrated in Fig. 6 for the same cell of two solenoids, one case with opposite polarity K=KK=-K and the other with the same polarity. In both cases, the propagation of beta functions is the same, whereas the effect on the phase advance is different; a splitting is observed. The solenoid cells were matched for $u=1/2$ and equal beta functions for a round beam. The propagation of phase advance in solenoid-type cells is given by
$\begin{split}\frac{d\mu_{1}}{ds}&=\frac{1-u}{\beta_{1x}}-\frac{R}{2}\sqrt{\frac{\beta_{1y}}{\beta_{1x}}}\sin\nu_{1},\\ \frac{d\mu_{2}}{ds}&=\frac{1-u}{\beta_{2y}}+\frac{R}{2}\sqrt{\frac{\beta_{2x}}{\beta_{2y}}}\sin\nu_{2}.\end{split}$ (43)
@@
Figure 6: Solenoid cells: Beta functions of opposite polarity of solenoid fields (top-left), same polarity (bottom-left). Phase advances of modes $1$ and 22 of opposite polarity (top-right), same polarity (bottom-right), for $u=1/2$.